Karen left her house heading east, and her sister left two hours later heading west at a rate 5 mph faster than Karen. Six hours after Karen left, she and her sister were 600 miles apart. How fast was Karen moving?
Answer by mananth(16949)

Karen left her house heading east,
her sister left two hours later heading west at a rate 5 mph faster than Karen.
Six hours after Karen left, she and her sister were 600 miles apart. How fast was Karen moving?
Karen speed be x mph
Karen has traveled 6x miles (in 6 hours)
..
sister x+5 mph .
Sister has traveled 4(x+5) miles. left 2 hours later.
..
4(x+5)+6x =600
4x+20+6x=600
10x+20=600
10x=600-20
10x=580
x= 58 mph speed of Karen